OpenStudy (anonymous):

Help? (3x^2 - 11x - 20)/(x^2 - 6x +5)

ganeshie8 (ganeshie8):

factor numerator factor denomoninator

ganeshie8 (ganeshie8):

and see if somethings cancel out

OpenStudy (anonymous):

I don't know how...

ganeshie8 (ganeshie8):

il help u factor numerator, you're going to factor denominator. deal ? :)

OpenStudy (anonymous):

Okay

ganeshie8 (ganeshie8):

3x^2 - 11x - 20 tell me two numbers such that :- 1) product = -20*3 = -60 2) sum = -11

OpenStudy (anonymous):

Uhh, 5 and 6 ? Idk

ganeshie8 (ganeshie8):

5*6 \(\ne\) -60

ganeshie8 (ganeshie8):

you need to look for factors of 60

ganeshie8 (ganeshie8):

-15*4 = -60 -15+4 = 11

ganeshie8 (ganeshie8):

so, -15 and 4 are our magic numbers

OpenStudy (anonymous):

4 and 15

ganeshie8 (ganeshie8):

yes, 4 and -15

ganeshie8 (ganeshie8):

3x^2 - 11x - 20 3x^2 - 15x + 4x - 20

ganeshie8 (ganeshie8):

i just wrote -11x as, -15x + 4x

ganeshie8 (ganeshie8):

fine ?

ganeshie8 (ganeshie8):

3x^2 - 11x - 20 3x^2 - 15x + 4x - 20 -------- ------- 3x(x-5) + 4(x-5)

OpenStudy (anonymous):

Oh

ganeshie8 (ganeshie8):

3x^2 - 11x - 20 3x^2 - 15x + 4x - 20 -------- ------- 3x(x-5) + 4(x-5) ---- ---- (x-5)(3x+4)

ganeshie8 (ganeshie8):

Done. let me knw if u really get my cryptic thing above :o

OpenStudy (anonymous):

I do. It's just hard for me since we just started this.

OpenStudy (anonymous):

For the denominator, after factoring, would it be x^2 - 3x + 2x +5 ?

ganeshie8 (ganeshie8):

x^2 - 6x +5 u need to find two numbers such that :- 1) product = 5*1 = 5 2) sum = -6

ganeshie8 (ganeshie8):

firs tell me, wat are the two magical numbers u picked

OpenStudy (anonymous):

Oooh, I did it wrong, I mixed them up. But it'd be -5 and 1

ganeshie8 (ganeshie8):

its -5 and -1

OpenStudy (anonymous):

Yeah.

ganeshie8 (ganeshie8):

try it again :)

OpenStudy (anonymous):

What

ganeshie8 (ganeshie8):

factor the denominator remember our deal ? :o

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

(x - 5)(5 - x) ?

ganeshie8 (ganeshie8):

step by step

OpenStudy (anonymous):

That's wrong..

ganeshie8 (ganeshie8):

x^2 - 6x +5 x^2 -1x - 5x + 5

ganeshie8 (ganeshie8):

i just wrote -6x as, -1x-5x

ganeshie8 (ganeshie8):

x^2 - 6x +5 x^2 -1x - 5x + 5 ------ ------

ganeshie8 (ganeshie8):

pull out GCF from first two terms pull out GCF from last two terms

OpenStudy (anonymous):

2 and 5 ?

ganeshie8 (ganeshie8):

first two terms, we can pull out oly x

ganeshie8 (ganeshie8):

x^2 - 6x +5 x^2 -1x - 5x + 5 ------ ------ x(x-1) - 5(x-1) ---- ---- (x-1)(x-5)

OpenStudy (anonymous):

Okay. I hope I can do the next one on my own

OpenStudy (anonymous):

Thanks :)

ganeshie8 (ganeshie8):

good :) factoring takes practice, just work 2-3 more problems. u wil feel confident im sure :)

OpenStudy (anonymous):

I'll try

ganeshie8 (ganeshie8):

good luck !